OpenStudy (anonymous):

Need help factoring. 16x^2-25

OpenStudy (anonymous):

difference of two square \[a^2-b^2=(a+b)(a-b)\]

OpenStudy (anonymous):

put \[a=4x,b=5\] to get \[(5x+4)(5x-4)\]

OpenStudy (saifoo.khan):

o.O

OpenStudy (callisto):

16x^2-25 = (4x)^2 - 5^2 = (4x-5)(4x+5)

OpenStudy (saifoo.khan):

Satellite are u sleeeeping this time???

OpenStudy (saifoo.khan):

(4x-5)(4x+5) tht's the right answer.

OpenStudy (anonymous):

zzzzzzzzzzzzzzzzzzzzzzzzzzzzzz

hero (hero):

lol

OpenStudy (anonymous):

Ok. Im still confused.

OpenStudy (anonymous):

i like to think of it as a typo

OpenStudy (saifoo.khan):

Lol.

OpenStudy (saifoo.khan):

yes u r right.

OpenStudy (anonymous):

\[(4x-5)(4x+5)\] multiply out, see that it is correct

OpenStudy (saifoo.khan):

thts right.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Thanks